OpenStudy (anonymous):

Find the soulution if possible 4p-6=8+2p

OpenStudy (anonymous):

p=6..y r uasking these?

OpenStudy (aroub):

4p-2p=6+8 2p=14 p=7

OpenStudy (anonymous):

there a grade and I dont understand them am taking up tutoring today for them there due in hr only have 3

OpenStudy (anonymous):

ok is it 6 or 7 pepps?

OpenStudy (anonymous):

peeps?

OpenStudy (aroub):

7